OpenStudy (anonymous):

Verify the identity. sin (x + pi/2) = cos x

hartnn (hartnn):

do you know a standard relation for sin (a+b) ?

hartnn (hartnn):

\(\sin (a+b)=\sin a \cos b +\cos a \sin b\)

hartnn (hartnn):

just plug in a =x and b =pi/2 there!

OpenStudy (anonymous):

what does cos(pi/2) and sin(pi/2) =?

hartnn (hartnn):

oh, its easy and standard cos (pi/2) = 0 sin (pi/2) = 1

OpenStudy (anonymous):

So I end up with sin x + cos x + sin (1)?

hartnn (hartnn):

no...show your work, i'll spot you the error

OpenStudy (anonymous):

sin(x)cos(pi/2) + cos(x)sin(pi/2) sin(x)cos(0) + cos(x)sin(1) cos(0) = 0 So, sin(x) + cos(x)sin(1) ?

hartnn (hartnn):

sin(x)cos(pi/2) + cos(x)sin(pi/2) cos pi/2 = 0 , sin pi/2 = 1 sin x (0) + cos x (1) =0+ cos x = cos x got it?!

hartnn (hartnn):

ask if any doubts...

OpenStudy (anonymous):

Oh!! I feel dumb now. I get it. Aha Thank you!! :)

hartnn (hartnn):

welcome ^_^